# RClicks: Realistic Click Simulation for Benchmarking Interactive Segmentation

### v1

@@ 1 Introduction
The task of interactive segmentation involves providing additional hints or prompts to the method, allowing it to produce more precise annotations compared to conventional semantic segmentation. The most famous member of interactive segmentation methods is Segment Anything (SAM) [kirillov2023segment, ravi2024sam]. Nowadays, SAM-like methods are applied in various fields, including the thin object segmentation [ke2023segment, xie2024pasam], medical segmentation [hu2023sam, zhou2023sam, biswas2023polypsam, zhang2023segment, SAM4MIS, mazurowski2023segment], 3D segmentation [yang2023sam3d, xu2023sampro3d], tracking [cheng2023segment] and video [ravi2024sam].
Typically, interactions occur in several rounds, where in each round the user corrects the prediction errors of the previous one. Evaluation of such methods requires user inputs. However, collecting many real-user inputs for multiple rounds is impractical since such a dataset needs to be rebuilt for every method and every interaction round due to its iterative nature. Thus, researchers often resort to a simple strategy to simulate user inputs. According to this strategy, a single click for each interaction round is generated as follows: (1) select the largest error region in the previous interaction round, and (2) click in the furthest point from the boundaries of this region (center point). Hereinafter, we refer to the above click sampling strategy as a baseline strategy, following [moskalenko2024tetris]. However, relying solely on this approach may result in overfitting and degraded performance in real-world usage scenarios.
(a) GrabCut [grabcut]
(b) Berkeley [berkeley-intro]
(c) COCO-MVal [coco]
(d) DAVIS [davis]
(e) TETRIS [moskalenko2024tetris]
Figure 1: Examples of real and predicted users’ clicks of interactive segmentation task. The upper row depicts real-users clicks (green) for a given target object (white contour); the middle and bottom rows visualize, correspondingly, clicks and their distribution predicted by our clickability model. Purple points in the middle and bottom rows represent clicks generated by the baseline strategy [ritm]. Mostly baseline click is close to a mode of users’ distribution (see 1(b) and 1(e)), however, in some cases it may be far from the mode (e.g. 1(a), 1(d)) or may not represent all modes of the distribution (e.g. 1(c), 1(e)).
@@ 2.1 User Input Types in Interactive Segmentation
Various types of user inputs have been explored in the literature. In [grabcut, ferrari2019scribbles] an initial selection is obtained using bounding boxes, and then refined with strokes. In [gueziri2017latency] object selection is done with strokes. [popenova2023contour] considers contours for selecting small objects, minor parts of an object, or a group of objects of the same type. [cheng2021flexible] proposes to use trimap, scribblemap or clickmap as an input. Segment Anything, or SAM [kirillov2023segment], processes multiple types of user prompts, including a point, a box, a mask, or a text.
Clicks-based approach selects objects of interest according to multiple user clicks (either positive or negative), and was first introduced in [deep-object-selection] and investigated in [ritm, liu2022simpleclick, sun2023cfricl, zhou2023interactivegpcis, kirillov2023segment, ke2023segment, zhang2023faster]. We focus solely on the click-based approach, since it is well-explored and has an established evaluation procedure in the field.
@@ 2.2 Benchmarking Interactive Segmentation
GrabCut [grabcut] is the first dataset proposed for interactive segmentation task. Then [berkeley-intro] adapted Berkeley [berkeley] segmentation dataset to evaluate interactive segmentation methods, but it required manual testing. However, manual testing is a time-consuming and resource-intensive process. Interactive segmentation expects multiple rounds of interactions, when each interaction depends on previous ones, and it is infeasible to apply manual procedure for larger scales. For these reasons, in practice, benchmarks generate user interactions automatically based on previous interactions.
In [deep-object-selection] authors proposed an automatic clicks generation strategy for evaluation on PASCAL VOC 2012 [pascal-voc-2012] and COCO [coco] segmentation datasets. The subsequent work [latent-diversity] used DAVIS [davis] and SBD [sbd] datasets for interactive segmentation, applying the same baseline strategy.
Most of the existing click-based methods [ritm, liu2022simpleclick, sun2023cfricl, zhou2023interactivegpcis, kirillov2023segment, ke2023segment, zhang2023faster] use the baseline strategy. However, it has not been validated in real-world usage scenarios until recently. [moskalenko2024tetris] introduced TETRIS benchmark and revealed that real users do not always click in the center of an area with the largest error, as assumed in the baseline strategy. Using the adversarial attacks, the paper demonstrated that methods have a tendency to overfit to the baseline strategy. Specifically, when the baseline clicks are used, the segmentation quality may be high, but even a slight change in the click position can result in a significant drop in quality. Therefore, the baseline strategy may not accurately estimate the quality of the methods in real usage. We believe that to estimate the actual quality, each click should be generated in accordance with human perception.
@@ 2.3 Saliency Prediction
The task of saliency prediction aims to model human perception by predicting probability maps [TranSalNet, kroner2020contextual] of user engagement in a free-view observation for a given media content. Reference data for this task usually comes from a specialized device – an eye tracker – which records eye fixations [judd2009learning, judd2012benchmark, CAT2000]. Subsequently, fixations from multiple viewers are aggregated into a probability distribution through Gaussian at each fixation point, with sigma corresponding to the retinal angle of a human’s field of view [judd2009learning]. Since scaling expensive eye tracker experiments is too complex, several researchers [jiang2015salicon, tavakoli2017saliency] proposed to use mouse movements as a proxy for saliency when training saliency models. However, saliency fixations cannot be directly used in the interactive segmentation task because saliency observers engage in free-viewing, while in our task, the user’s goal is to make a click to highlight a specific object or a part of it. Thus, for the interactive segmentation problem, real-user clicks should be collected.
@@ 3 Users’ Clicks Dataset
We propose a novel dataset of real-users clicks for interactive segmentation. Our dataset is based on the existing image segmentation datasets. In total, we collected 475 544 user inputs for GrabCut [grabcut], Berkeley [berkeley], DAVIS [davis], COCO-MVal [lin2014microsoft], TETRIS [moskalenko2024tetris]. To gather users’ clicks, we developed a specialized presentation tool. Specifically, in each task, we asked users to click on the target objects by displaying images and corresponding segmentation masks. We considered several display modes to instruct users what objects should be selected. As interface can cause bias in clicks distribution, we conducted a user study to select the option that best mimics natural user object selections.
@@ 3.1 Collection Procedure
We considered the following task displaying modes. (a) Text Descriptionmode shows the textual description of the target object for 2.5 seconds. (b) Object CutOutvisualizes the target instance in its original position on a gray background for 2 seconds. (c) During Shifted CutOut, the target instance is shown for 2 seconds on a gray background, then shifted to the top-left corner, which aims to motivate the assessor to independently locate the instance on the image as its position is shifted. (d) Silhouette Maskshows a black-and-white mask of the target instance in the original position for 2 seconds. (e) Highlighted Instancedisplays the original image with the background where the target instance is highlighted with a green border. Rationale for our choice of these displaying modes and time periods can be found in Appendix B.1.
@@ 3.2 Selecting Unbiased Task Display Mode
Figure 3: The difficulty of utilizing text descriptions when instructing participants on the first round selection of a certain instance (3(a)-3(b)) and to select or unselect the certain error area in the subsequent round (3(c)-3(d)).
Therefore, we conducted an ablation study on 100 randomly selected images from TETRIS dataset [moskalenko2024tetris]. We used images and segmentation masks from TETRIS, and additionally manually annotated textual descriptions. In this study, we compare the clicks obtained via considered display modes with clicks collected through Text Description.
@@
To compare quantitatively display modes with unbiased Text Description mode, we utilized the following sample-based metrics: (a) PL1 – mean of all pairwise L1-distances between click coordinates, normalized by object width and height. (b) WD – Wasserstein distance [peyre2020computational] between click coordinates, normalized by object width and height. (c) KS – Kolmogorov-Smirnov test in 2D case [christian2020re, peacock1983two, fasano1987multidimensional]. We conclude that clicks are not significantly different if a p-value is greater than 0.05. The indicator function is used as a metric, which equals 1 when clicks are not significantly different.
@@ 3.3 Collected Interactions
We annotated each instance in all common interactive segmentation benchmark datasets – DAVIS [davis], GrabCut [grabcut], COCO-MVal [coco], Berkeley [berkeley], TETRIS [moskalenko2024tetris] using PC and mobile clicks. Collected clicks were validated similarly to the ablation stage. When annotating subsequent interaction rounds, the user should click in the area of the segmentation error. To obtain error masks for the subsequent rounds, we applied state-of-the-art interactive segmentation methods – SAM [kirillov2023segment], SimpleClick [liu2022simpleclick], and RITM [ritm] – to all images and all clicks corresponding to those images from the first round. Then, for each image, we selected the mask with the highest quality up to a threshold of 0.95 IoU. We motivate it by the fact that at such a high level of quality, the user is likely to stop annotating the instance as the errors would be minimal, and even the radius of the click may exceed the size of the erroneous area. In total, we collected 475 thousand valid clicks from users. The number of valid clicks annotated for each dataset and interaction round is presented in Table 2.
@@ 4.1 Baseline Models
As baselines for comparison, we considered uniform, distance, and saliency distribution models. The uniform hypothesis postulates that the clickability of all pixels within the target area is equally distributed (see Figure 4(b)). When the area of interest is relatively small, the uniform assumption is reasonable. However, according to this assumption, the click probability of object boundaries and their centers is equal, which is not necessarily true. The distance transform [moskalenko2024tetris, ritm] addresses this issue by assigning greater weight to pixels in the center of the object than to those on the boundary (see Figure 4(c)). Nevertheless, this transform considers only the shape of the object, neglecting human perception. To account for human perception, saliency distribution can be used. This is a reasonable baseline, as users look at the target area of interaction when clicking. For the saliency baseline, we utilized a state-of-the-art model, TranSalNet [lou2022transalnet]. The example of constructed saliency distribution is presented in Figure 4(d), details of how such map is constructed can be found in the Appendix A.2. However, saliency models are trained for free-viewing task, and do not take into consideration the setup of our task.
@@ 4.2 Clickability Prediction Model
Similar to the saliency prediction task, we formulate the task of simulating user clicks as a probabilistic problem. Given an image, a ground truth object mask, and a segmentation error mask (FP or FN), the model should predict at each pixel the probability of being clicked. We refer to this as a clickability map (see details in Section 4.3). The proposed pipeline is shown in Figure 5. As a base architecture for our model, we adapted state-of-the-art SegNeXt segmentation network [guo2022segnext]. We input the original image into the network and concatenate the ground truth mask with the error mask, feeding the resulting tensor as an additional input to the network, a technique inspired by the Conv1S [ritm]. We use the Kullback-Leibler divergence (KLD) loss function between the predicted and ground truth distributions.
@@ 4.3 Clickability Maps Dataset
We constructed train and validation datasets as follows. We split images of TETRIS dataset into non-overlapping train and validation parts. Since we do not know the real click density, for model training and validation, several sets of clickability maps were constructed with varying magnitudes of sigma from step (3). To choose the best sigma, we conducted an ablation study, which can be found in Appendix A.3. Note that we constructed clickability maps using clicks from both smartphones and PCs. This was done to ensure that the model would predict clicks regardless of the device type.
@@ 4.4 Models Evaluation
To choose the best clickability model, we evaluated considered models on the real-user clicks of TETRIS validation part. Here, in addition to sample-based metrics considered above, we calculated additional metrics, that were computed based on the ground-truth clicks positions and predicted clicks distribution: PDE – likelihood of ground-truth clicks, and NSS from saliency benchmarks [kummerer2018saliency]. Evaluation results are presented in Table 3. Our clickability model shows the best performance. We address the question of model generalizability in the Appendix A.2.
@@ 5 Benchmarking Interactive Segmentation
We modify a common evaluation protocol [ritm] by replacing the baseline sampling strategy with sampling from different clicking groups, obtained through the clickability model. Specifically, (1) interactive segmentation metrics (e.g. NoC) are calculated for every instance in a dataset and group Gi by sampling click from Gi (weighted by the clickability model) for every interaction round (in our experiments – 20 rounds); (2) then for each instance statistics (e.g. mean and standard deviation) of sampled metrics over clicking groups are estimated; (3) finally, these statistics are averaged over all instances in the dataset.
@@
In addition to the evaluation results on simulated clicks, we provide evaluation results on the first round real clicks. The evaluation on the subsequent real clicks is infeasible, since interactive segmentation in a subsequent round depends on a model output from a previous round. However, there is no such problem in the first round, and actual performance metrics can be computed on the real clicks of the first round. Therefore, we employed the real clicks from the first round as follows: (1) computed real-world accuracy (see Appendix C); (2) compared accuracy of interactive segmentation methods using real and simulated clicks (see Appendix C); and (3) estimated real-world robustness of the methods for each instance in the dataset. We estimated the latter through IoU noise-to-signal ratio (NSR). The greater the value of NSR, the more difficult such an instance is for segmentation. Figure 8 plots a scatter of mean vs. standard deviation of IoU over real-users first clicks and segmentation methods from Table 4.
@@
Method Backbone Data DAVIS [davis] COCO-MVal [coco] TETRIS [moskalenko2024tetris]
@@
GPCIS [zhou2023interactivegpcis] RN50 C+L 6.44±0.85 16.88 53.65 4.74±1.31 26.43 79.00 3.87±0.79 19.55 56.43
CDNet [chen2021conditional] RN34 C+L 5.95±0.73 14.95 45.88 4.13±0.85 15.15 49.79 3.10±0.53 14.16 44.06
@@
CFR-ICL [sun2023cfricl] ViT-H C+L 4.53±0.46 9.32 18.47 2.70±0.63 9.58 24.13 2.12±0.34 8.76 14.33
MobileSAM [zhang2023faster] ViT-Tiny SA-1B 5.96±0.56 8.63 15.39 5.25±0.78 9.79 19.78 3.42±0.48 7.47 12.69
SAM [kirillov2023segment] ViT-B SA-1B 5.30±0.53 8.26 11.27 4.91±0.79 9.88 15.73 3.04±0.51 11.17 10.06

### v2

@@ 1 Introduction
The task of interactive segmentation involves providing additional hints or prompts to the method, allowing it to produce more precise annotations compared to conventional semantic segmentation. The most famous member of interactive segmentation methods is Segment Anything (SAM) [1, 2]. Nowadays, SAM-like methods are applied in various fields, including the thin object segmentation [3, 4], medical segmentation [5, 6, 7, 8, 9, 10], 3D segmentation [11, 12], tracking [13] and video [2].
Typically, interactions occur in several rounds, where in each round the user corrects the prediction errors of the previous one. Evaluation of such methods requires user inputs. However, collecting many real-user inputs for multiple rounds is impractical since such a dataset needs to be rebuilt for every method and every interaction round due to its iterative nature. Thus, researchers often resort to a simple strategy to simulate user inputs. According to this strategy, a single click for each interaction round is generated as follows: (1) select the largest error region in the previous interaction round, and (2) click in the furthest point from the boundaries of this region (center point). Hereinafter, we refer to the above click sampling strategy as a baseline strategy, following [14]. However, relying solely on this approach may result in overfitting and degraded performance in real-world usage scenarios.
(a) GrabCut [15]
(b) Berkeley [16]
(c) COCO-MVal [17]
(d) DAVIS [18]
(e) TETRIS [14]
Figure 1: Examples of real and predicted users’ clicks of interactive segmentation task. The upper row depicts real-users clicks (green) for a given target object (white contour); the middle and bottom rows visualize, correspondingly, clicks and their distribution predicted by our clickability model. Purple points in the middle and bottom rows represent clicks generated by the baseline strategy [19]. Mostly baseline click is close to a mode of users’ distribution (see 1(b) and 1(e)), however, in some cases it may be far from the mode (e.g. 1(a), 1(d)) or may not represent all modes of the distribution (e.g. 1(c), 1(e)).
@@ 2.1 User Input Types in Interactive Segmentation
Various types of user inputs have been explored in the literature. In [15, 20] an initial selection is obtained using bounding boxes, and then refined with strokes. In [21] object selection is done with strokes. [22] considers contours for selecting small objects, minor parts of an object, or a group of objects of the same type. [23] proposes to use trimap, scribblemap or clickmap as an input. Segment Anything, or SAM [1], processes multiple types of user prompts, including a point, a box, a mask, or a text.
Clicks-based approach selects objects of interest according to multiple user clicks (either positive or negative), and was first introduced in [24] and investigated in [19, 25, 26, 27, 1, 3, 28]. We focus solely on the click-based approach, since it is well-explored and has an established evaluation procedure in the field.
@@ 2.2 Benchmarking Interactive Segmentation
GrabCut [15] is the first dataset proposed for interactive segmentation task. Then [16] adapted Berkeley [29] segmentation dataset to evaluate interactive segmentation methods, but it required manual testing. However, manual testing is a time-consuming and resource-intensive process. Interactive segmentation expects multiple rounds of interactions, when each interaction depends on previous ones, and it is infeasible to apply manual procedure for larger scales. For these reasons, in practice, benchmarks generate user interactions automatically based on previous interactions.
In [24] authors proposed an automatic clicks generation strategy for evaluation on PASCAL VOC 2012 [30] and COCO [17] segmentation datasets. The subsequent work [31] used DAVIS [18] and SBD [32] datasets for interactive segmentation, applying the same baseline strategy.
Most of the existing click-based methods [19, 25, 26, 27, 1, 3, 28] use the baseline strategy. However, it has not been validated in real-world usage scenarios until recently. [14] introduced TETRIS benchmark and revealed that real users do not always click in the center of an area with the largest error, as assumed in the baseline strategy. Using the adversarial attacks, the paper demonstrated that methods have a tendency to overfit to the baseline strategy. Specifically, when the baseline clicks are used, the segmentation quality may be high, but even a slight change in the click position can result in a significant drop in quality. Therefore, the baseline strategy may not accurately estimate the quality of the methods in real usage. We believe that to estimate the actual quality, each click should be generated in accordance with human perception.
@@ 2.3 Saliency Prediction
The task of saliency prediction aims to model human perception by predicting probability maps [33, 34] of user engagement in a free-view observation for a given media content. Reference data for this task usually comes from a specialized device – an eye tracker – which records eye fixations [35, 36, 37]. Subsequently, fixations from multiple viewers are aggregated into a probability distribution through Gaussian at each fixation point, with sigma corresponding to the retinal angle of a human’s field of view [35]. Since scaling expensive eye tracker experiments is too complex, several researchers [38, 39] proposed to use mouse movements as a proxy for saliency when training saliency models. However, saliency fixations cannot be directly used in the interactive segmentation task because saliency observers engage in free-viewing, while in our task, the user’s goal is to make a click to highlight a specific object or a part of it. Thus, for the interactive segmentation problem, real-user clicks should be collected.
@@ 3 Users’ Clicks Dataset
We propose a novel dataset of real-users clicks for interactive segmentation. Our dataset is based on the existing image segmentation datasets. In total, we collected 475 544 user inputs for GrabCut [15], Berkeley [29], DAVIS [18], COCO-MVal [40], TETRIS [14]. To gather users’ clicks, we developed a specialized presentation tool. Specifically, in each task, we asked users to click on the target objects by displaying images and corresponding segmentation masks. We considered several display modes to instruct users what objects should be selected. As interface can cause bias in clicks distribution, we conducted a user study to select the option that best mimics natural user object selections.
@@ 3.1 Collection Procedure
We considered the following task displaying modes. (a) Text Descriptionmode shows the textual description of the target object for 2.5 seconds. (b) Object CutOutvisualizes the target instance in its original position on a gray background for 2 seconds. (c) During Shifted CutOut, the target instance is shown for 2 seconds on a gray background, then shifted to the top-left corner, which aims to motivate the assessor to independently locate the instance on the image as its position is shifted. (d) Silhouette Maskshows a black-and-white mask of the target instance in the original position for 2 seconds. (e) Highlighted Instancedisplays the original image with the background where the target instance is highlighted with a green border. Rationale for our choice of these displaying modes and time periods can be found in Appendix C.1.
@@ 3.2 Selecting Unbiased Task Display Mode
Figure 3: Examples of situations where instructing participants with text descriptions may be challenging or ambiguous: selection of a certain instance in the first round (3(a)-3(b)); and selecting or unselecting a certain error area in the subsequent round (3(c)-3(d)).
Therefore, we conducted an ablation study on 100 randomly selected images from TETRIS dataset [14]. We used images and segmentation masks from TETRIS, and additionally manually annotated textual descriptions. In this study, we compare the clicks obtained via considered display modes with clicks collected through Text Description.
@@
To compare quantitatively display modes with unbiased Text Description mode, we utilized the following sample-based metrics: (a) PL1 – mean of all pairwise L1-distances between click coordinates, normalized by object width and height. (b) WD – Wasserstein distance [41] between click coordinates, normalized by object width and height. (c) KS – Kolmogorov-Smirnov test in 2D case [42, 43, 44]. We conclude that clicks are not significantly different if a p-value is greater than 0.05. The indicator function is used as a metric, which equals 1 when clicks are not significantly different.
@@ 3.3 Collected Interactions
We annotated each instance in all common interactive segmentation benchmark datasets – DAVIS [18], GrabCut [15], COCO-MVal [17], Berkeley [29], TETRIS [14] using PC and mobile clicks. Collected clicks were validated similarly to the ablation stage. When annotating subsequent interaction rounds, the user should click in the area of the segmentation error. To obtain error masks for the subsequent rounds, we applied state-of-the-art interactive segmentation methods – SAM [1], SimpleClick [25], and RITM [19] – to all images and all clicks corresponding to those images from the first round. Then, for each image, we selected the mask with the highest quality up to a threshold of 0.95 IoU. We motivate it by the fact that at such a high level of quality, the user is likely to stop annotating the instance as the errors would be minimal, and even the radius of the click may exceed the size of the erroneous area. In total, we collected 475 thousand valid clicks from users. The number of valid clicks annotated for each dataset and interaction round is presented in Table 2.
@@ 4.1 Baseline Models
As baselines for comparison, we considered uniform, distance, and saliency distribution models. The uniform hypothesis postulates that the clickability of all pixels within the target area is equally distributed (see Figure 4(b)). When the area of interest is relatively small, the uniform assumption is reasonable. However, according to this assumption, the click probability of object boundaries and their centers is equal, which is not necessarily true. The distance transform [14, 19] addresses this issue by assigning greater weight to pixels in the center of the object than to those on the boundary (see Figure 4(c)). Nevertheless, this transform considers only the shape of the object, neglecting human perception. To account for human perception, saliency distribution can be used. This is a reasonable baseline, as users look at the target area of interaction when clicking. For the saliency baseline, we utilized a state-of-the-art model, TranSalNet [45]. The example of constructed saliency distribution is presented in Figure 4(d), details of how such map is constructed can be found in the Appendix B.2. However, saliency models are trained for free-viewing task, and do not take into consideration the setup of our task.
@@ 4.2 Clickability Prediction Model
Similar to the saliency prediction task, we formulate the task of simulating user clicks as a probabilistic problem. Given an image, a ground truth object mask, and a segmentation error mask (FP or FN), the model should predict at each pixel the probability of being clicked. We refer to this as a clickability map (see details in Section 4.3). The proposed pipeline is shown in Figure 5. As a base architecture for our model, we adapted state-of-the-art SegNeXt segmentation network [46]. We input the original image into the network and concatenate the ground truth mask with the error mask, feeding the resulting tensor as an additional input to the network, a technique inspired by the Conv1S [19]. We use the Kullback-Leibler divergence (KLD) loss function between the predicted and ground truth distributions.
@@ 4.3 Clickability Maps Dataset
We constructed train and validation datasets as follows. We split images of TETRIS dataset into non-overlapping train and validation parts. Since we do not know the real click density, for model training and validation, several sets of clickability maps were constructed with varying magnitudes of sigma from step (3). To choose the best sigma, we conducted an ablation study, which can be found in Appendix B.3. Note that we constructed clickability maps using clicks from both smartphones and PCs. This was done to ensure that the model would predict clicks regardless of the device type.
@@ 4.4 Models Evaluation
To choose the best clickability model, we evaluated considered models on the real-user clicks of TETRIS validation part. Here, in addition to sample-based metrics considered above, we calculated additional metrics, that were computed based on the ground-truth clicks positions and predicted clicks distribution: PDE – likelihood of ground-truth clicks, and NSS from saliency benchmarks [47]. Evaluation results are presented in Table 3. Our clickability model shows the best performance. We address the question of model generalizability in the Appendix B.2.
@@ 5 Benchmarking Interactive Segmentation
We modify a common evaluation protocol [19] by replacing the baseline sampling strategy with sampling from different clicking groups, obtained through the clickability model. Specifically, (1) interactive segmentation metrics (e.g. NoC) are calculated for every instance in a dataset and group Gi by sampling click from Gi (weighted by the clickability model) for every interaction round (in our experiments – 20 rounds); (2) then for each instance statistics (e.g. mean and standard deviation) of sampled metrics over clicking groups are estimated; (3) finally, these statistics are averaged over all instances in the dataset.
@@
In addition to the evaluation results on simulated clicks, we provide evaluation results on the first round real clicks. The evaluation on the subsequent real clicks is infeasible, since interactive segmentation in a subsequent round depends on a model output from a previous round. However, there is no such problem in the first round, and actual performance metrics can be computed on the real clicks of the first round. Therefore, we employed the real clicks from the first round as follows: (1) computed real-world accuracy (see Appendix D); (2) compared accuracy of interactive segmentation methods using real and simulated clicks (see Appendix D); and (3) estimated real-world robustness of the methods for each instance in the dataset. We estimated the latter through IoU noise-to-signal ratio (NSR). The greater the value of NSR, the more difficult such an instance is for segmentation. Figure 8 plots a scatter of mean vs. standard deviation of IoU over real-users first clicks and segmentation methods from Table 4.
@@
Method Backbone Data DAVIS [18] COCO-MVal [17] TETRIS [14]
@@
GPCIS [27] RN50 C+L 6.44±0.85 16.88 53.65 4.74±1.31 26.43 79.00 3.87±0.79 19.55 56.43
CDNet [48] RN34 C+L 5.95±0.73 14.95 45.88 4.13±0.85 15.15 49.79 3.10±0.53 14.16 44.06
@@
CFR-ICL [26] ViT-H C+L 4.53±0.46 9.32 18.47 2.70±0.63 9.58 24.13 2.12±0.34 8.76 14.33
MobileSAM [28] ViT-Tiny SA-1B 5.96±0.56 8.63 15.39 5.25±0.78 9.79 19.78 3.42±0.48 7.47 12.69
SAM [1] ViT-B SA-1B 5.30±0.53 8.26 11.27 4.91±0.79 9.88 15.73 3.04±0.51 11.17 10.06